A compound contains 28 percent foo nitrogen and 72 percent of metal by weight. 3 atom of metal combine with 2 atoms find the atomic weight of metal.              

Grade:10

2 Answers

Jitender Pal
9 years ago
According to problem, three atoms of M combine with 2 atoms of N
Equivalent wt of N = 14/3 (∵ valency of N in compound is 3)
∵ 288 g n combines with = 72g metal
∴ 14/3 N combines with = 72/28 * 14/3 = 12
∴ Eq. wt. of metal = 12
At wt of metal = Eq. wt * valency = 12 * 2 = 24 [Valency]
ishan jogalekar
6 years ago
100-28 =72% of metal is there in the element also the mass given is 28 g of N.sosame 72 g of metal is presenting compound as it`s molecular wt.but as there are 3 atoms of metal present it will be 72÷3=24